OpenStudy (anonymous):

A girl is now one-third as old as her mother. In three years, she will be two-fifths as old as her mother will be. What are their present ages?

OpenStudy (zzr0ck3r):

g=(1/3)m (g+5) = (2/5)(m+5) these are our two equations 3g = m 5g+25 = 2m+10 5g+25 = 2(3g)+10 5g+25=6g+10 g = 15 thus m = 45